# The Role of Graph Topology in the Performance of Biomedical Knowledge Graph Completion Models

### v1

@@ Abstract
Knowledge Graph Completion has been increasingly adopted as a useful method for several tasks in biomedical research, like drug repurposing or drug-target identification. To that end, a variety of datasets and Knowledge Graph Embedding models has been proposed over the years. However, little is known about the properties that render a dataset useful for a given task and, even though theoretical properties of Knowledge Graph Embedding models are well understood, their practical utility in this field remains controversial. We conduct a comprehensive investigation into the topological properties of publicly available biomedical Knowledge Graphs and establish links to the accuracy observed in real-world applications. By releasing all model predictions and a new suite of analysis tools we invite the community to build upon our work and continue improving the understanding of these crucial applications.
Machine Learning, ICML, Knowledge Graphs, Knowledge Graph Completion, Graph Topology, Drug Discovery, Biomedical Research
@@ 1 Introduction
Knowledge Graphs (KGs) have become a powerful way of integrating information about a domain into a single structured representation. KGs represent knowledge as a collection of facts in the form of triples $(h,r,t)$, which make a statement about the relationship $r$ between two entities $h$ and $t$. In KGs from the biomedical domain, entities are genes, diseases, drugs, or pathways (among others), whilst relations represent biological interactions or associations between them (MacLean, 2021; Bonner et al., 2022a; Zeng et al., 2022). Biomedical KGs have been used to support a variety of tasks within the complex and costly drug discovery process, such as drug repurposing (Bang et al., 2023), identifying synergistic drug combinations (Zitnik et al., 2018) and target discovery (Paliwal et al., 2020). Unlike in other domains, biomedical KGs frequently comprise information from a wide range of conceptual abstraction levels. Edges representing information with a low abstraction level, such as experimentally validated protein-protein interactions (Protein A, interacts, Protein B), are found in the same KG as mid abstraction level edges, such as associations between genes and diseases taken from a genome-wide association study (Gene A, associates, Disease X), and high abstraction level edges, such as human-curated ontology hierarchies (Disease X, subclass of, Disease Y). This makes biomedical KGs uniquely interesting to consider when they are used in combination with Machine Learning algorithms, as models must learn to make predictions across these different abstraction levels simultaneously.
@@
Since KGs, including biomedical KGs, are typically incomplete, inferring missing triples is a key Machine Learning application – a task known as Knowledge Graph Completion in the literature. Perhaps the most common method of making predictions with KGs involves using Knowledge Graph Embedding (KGE) models, which learn low dimensional representations for all entities and relation types in a KG and combine them through a scoring function to compute the likelihood of a missing edge. In the biomedical domain, Knowledge Graph Completion has been used for a diverse set of tasks including predicting new links between genes and diseases (Paliwal et al., 2020). However, little is known about the practical ability of ML models to leverage topological patterns in the graph in order to make better predictions. An example of such a pattern is composition, where a relation between two entities is inferred by the presence of a shared third entity. In the biomedical domain, this could be the case when a drug is inferred to treat a disease because they share a common gene connection. Ideally, we would like to discover new drug repurposing opportunities by learning from such patterns in the data. Similarly, the inference pattern uses the presence of a certain relation type between an entity pair to infer the presence of a different relation between the same pair, which could be used to discover novel therapeutic gene-disease connections by aggregating multiple evidence types.
@@
While the theoretical implications of these topological patterns for KGE models are well understood – in fact, new KGE models are often designed to capture key topological patterns – their impact on the predictive performance in real-world Knowledge Graph Completion tasks remains an open question. Recent work has started to investigate this (Ali et al., 2021; Jin et al., 2023; Teneva & Hruschka, 2023), but has primarily focussed on the properties of relation types rather than individual triples. When comparing or proposing new KGE models, this approach helps to understand which topological patterns can be captured by the relation embedding. However, when investigating the impact of the graph structure on KGE model performance, this relation level approach might neglect critical micro-level structural data. Moreover, existing work has not focussed specifically on the biomedical domain, where there is a clear need for understanding the relationship between KG topology and model performance. New KGE models are typically developed and evaluated on a small set of general-domain KGs, which may contain distinct topological patterns compared to those from the biomedical domain. Biomedical KGs are often constructed by practitioners to address a specific purpose, requiring decisions about the key entities and relation types to include, the optimal data sources from which to extract information and how best to model the desired knowledge in a KG schema. This manual graph composition process is costly and time-consuming, an issue that can be compounded when KGs are designed as input for ML models, often requiring several iterations to best tune the graph to the predictive task at hand (Bonner et al., 2022a; Ratajczak et al., 2022). Thus, there is a growing interest in understanding how any decisions taken which affect topological properties of a KG can ultimately impact predictions made from them.
In this paper, we explore the relationship between biomedical KG topology and KGE model accuracy. Our main contributions are as follows:
@@
  • We provide an in-depth analysis of the topological properties of six public KGs, focussing on the biomedical domain, and compare the corresponding predictive performance of four well-established KGE models. While previous work has limited the investigation at the macro-level of relation types, we are able to detect stronger patterns linking topological properties to predictive accuracy by zooming in at the level of individual triples.
@@
  • To aid future work in this area, we propose a standardized framework to describe KG topology and release a dedicated Python package kg-topology-toolbox implementing it [^1] https://github.com/graphcore-research/kg-topology-toolbox.
Finally, to help bridge the gap between industry and academy, we make the predictions of the trained models available to the community (Cattaneo et al., 2024) to conduct further analysis.
@@ 2 Related Work
Since the adoption of representation learning for Knowledge Graph Completion, there has been interest in linking predictive performance of models to topological structures in the underlying Knowledge Graphs (Bordes et al., 2013). These structures, taking into account the heterogeneous relation types in KGs, have been split into two categories: edge cardinalities and topological patterns (see Section 3 for the definitions used in this study). Indeed, it has become common for new models to be designed with the capacity of capturing these patterns (Cao et al., 2021; Chao et al., 2021; Sun et al., 2019), working under the assumption that better modelling them will result in improved predictive performance (Jin et al., 2023).
@@
There has been limited work providing a comprehensive study on the impact of these patterns across different models and datasets, although some studies have started to appear recently. For example, an investigation across 3 common benchmark KGs (FB15k-237, WN18RR and Yago310) and 21 KGE models found that symmetric relations were easier to predict that anti-symmetric and compositions ones (Ali et al., 2021). Another study compared the performance of 7 KGE models on FB15k-237 and WN18RR, focussing on 4 different topological patterns (Jin et al., 2023), finding that a model’s theoretical capacity to capture a certain pattern does not guarantee good predictive performance on it. However, none of these studies explores the impact of edge patterns on biomedical KGs specifically.
KGE models have been shown to be strongly impacted by entity degree, with models often being heavily biased towards highly connected entities (Bonner et al., 2022b; Rossi et al., 2021b, a; Mohamed et al., 2020). However, disentangling the effect of degree from other topological properties when considering model performance is often overlooked by prior studies. One study showed that as entity degree increases, performance on symmetric relations decreases (Jin et al., 2023). However, it is still an open research question as to how degree obscures other correlations in the performance of models on different topological patterns.
The prevalence at which different topological patterns occur in real-world datasets is also an active area of consideration. Teneva & Hruschka (2023) investigated the frequency of topological and cardinality patterns in 29 real-world KGs from a range of domains, including biomedical. They observed that, under their chosen pattern definitions, many relations in biomedical KGs are predominantly many-to-many and that there was a lack of composition relations. Other work investigated the prevalence of symmetric and composition relations in 3 common KG benchmark datasets, finding that the majority of relations were anti-symmetric and that composition relations were rarer (Ali et al., 2021).
@@ 3 Knowledge Graph Topological Properties
For a triple $(h,r,t)$ in a KG $\mathcal{G}$ we define the head out-degree as the cardinality of the set $\{\hat{t}\mid\exists\hat{r}:(h,\hat{r},\hat{t})\in\mathcal{G}\}$ and, analogously, the tail in-degree as the cardinality of $\{\hat{h}\mid\exists\hat{r}:(\hat{h},\hat{r},t)\in\mathcal{G}\}$. We further define the head out-degree of same relation $\textrm{deg}_{r}(h)$ and the tail in-degree of same relation $\textrm{deg}_{r}(t)$ as the cardinalities of $\{\hat{t}\mid(h,r,\hat{t})\in\mathcal{G}\}$ and $\{\hat{h}\mid(\hat{h},r,t)\in\mathcal{G}\}$, respectively.
The edge cardinality of a triple $(h,r,t)$ is then defined as follows:
$\textrm{deg}_{r}(h)$ $\textrm{deg}_{r}(t)$ $=1$ $>1$
$=1$ one-to-one many-to-one
$>1$ one-to-many many-to-many
@@
Figure 1: Example of edge cardinalities in a KG with two relation types (blue, red).
@@
[…]
(a) Symmetry
[…]
(b) Inference
[…]
(c) Inverse
[…]
(d) Composition
Figure 2: Edge topological patterns. We refer to the dashed edge(s) as counterpart edge of $(h,r,t)$.
@@
While the above cardinalities and topological patterns have an intuitive definition for individual triples, previous studies (Jin et al., 2023; Teneva & Hruschka, 2023; Bordes et al., 2013; Sun et al., 2019) have considered them as properties of relation types as a whole (e.g., relation $r$ is said to be symmetric if $\>\forall h,t,\;\displaystyle{(h,r,t)\Rightarrow(t,r,h)}$), despite the fact that a given property might be satisfied only by a fraction of the triples of a certain relation. We find that such conceptual generalization lacks a proper formalization in the literature, with inconsistencies in how it is performed, and is prone to introducing noise in results (see Section 5.2). Thus, we consider the above defined properties only as properties of triples.
@@ 4 Experimental Setup
We investigate five public biomedical KGs: Hetionet (Himmelstein et al., 2017), PrimeKG [^2] In the original PrimeKG graph, for every triple $(h,r,t)$ the reverse triple $(t,r,h)$ is present as well. We prepocessed PrimeKG to remove these reverse edges. (Chandak et al., 2023), PharmKG (Zheng et al., 2021), OpenBioLink (Breit et al., 2020), PharMeBINet (Königs et al., 2022) (detailed in Table A.1). We also include the trivia KG FB15k-237 (Toutanova & Chen, 2015) as a baseline, to detect any results unique to the biomedical domain. On all KGs we train four of the most popular KGE models: TransE (Bordes et al., 2013), DistMult (Yang et al., 2015), RotatE (Sun et al., 2019) and TripleRE (Yu et al., 2022). Whilst being all well-established among practitioners (Hu et al., 2020), these models differ in complexity and in the relational patterns they can capture (Table A.2). In particular, TransE and DistMult model the interaction between head entity, relation type and tail entity in quite simple terms. Nevertheless, despite missing certain topological patterns, they remain strong baselines. On the other hand, RotatE and TripleRE are more recent approaches able to capture all four investigated topological patterns. Notice however that the theoretical capability of a scoring function to model a particular edge topological pattern is not per se a guarantee of stronger predictive performance on such edges (Jin et al., 2023). Our experiments are designed to quantify this impact.

### v2

@@ Abstract
Knowledge Graph Completion has been increasingly adopted as a useful method for helping address several tasks in biomedical research, such as drug repurposing or drug-target identification. To that end, a variety of datasets and Knowledge Graph Embedding models have been proposed over the years. However, little is known about the properties that render a dataset, and associated modelling choices, useful for a given task. Moreover, even though theoretical properties of Knowledge Graph Embedding models are well understood, their practical utility in this field remains controversial. In this work, we conduct a comprehensive investigation into the topological properties of publicly available biomedical Knowledge Graphs and establish links to the accuracy observed in real-world tasks. By releasing all model predictions and a new suite of analysis tools we invite the community to build upon our work and continue improving the understanding of these crucial applications.
Machine Learning, Knowledge Graphs, Knowledge Graph Completion, Graph Topology, Drug Discovery, Biomedical Research
@@ 1 Introduction
Knowledge Graphs (KGs) have become a powerful way of integrating information about a domain into a single structured representation. KGs represent knowledge as a collection of facts in the form of triples $(h,r,t)$, which make a statement about the relationship $r$ between two entities $h$ and $t$. In KGs from the biomedical domain, entities can be genes, diseases, drugs, or pathways (among others), whilst relations represent biological interactions or associations between them (MacLean, 2021; Bonner et al., 2022a; Zeng et al., 2022). Biomedical KGs have been used to support a variety of tasks within the complex and costly drug discovery process, such as drug repurposing (Bang et al., 2023), identifying synergistic drug combinations (Zitnik et al., 2018) and target discovery (Paliwal et al., 2020). Unlike in other domains, biomedical KGs frequently comprise information from a wide range of conceptual abstraction levels, for both entities and edges. It is common to find edges representing information with a low abstraction level, such as experimentally validated protein-protein interactions (Protein A, interacts, Protein B), in the same KG as mid abstraction level edges, such as associations between genes and bioprocesses (Gene A, associates, BioProcess X), and high abstraction level edges, such as disease ontology hierarchies (Disease X, subclass of, Disease Y). Entities can also have different abstraction levels, with some representing physical entities (e.g., genes) and others representing conceptual entities (e.g., diseases or biological processes). This makes biomedical KGs uniquely interesting to consider when they are used in combination with Machine Learning algorithms, as models must learn to make predictions across these different abstraction levels simultaneously.
@@
Since KGs, including biomedical ones, are typically incomplete, inferring missing triples is a key Machine Learning application – a task known as Knowledge Graph Completion in the literature. Perhaps the most common method of making predictions on KGs involves using Knowledge Graph Embedding (KGE) models, which learn low dimensional representations for all entities and relation types in a KG and combine them through a scoring function to compute the likelihood of a missing edge. In the biomedical domain, Knowledge Graph Completion has been used for a diverse set of tasks including predicting new links between genes and diseases (Paliwal et al., 2020) and possible drug interactions, among others (Gema et al., 2024).
Biomedical KGs are often constructed by practitioners to address a specific purpose (e.g., investigate the pathways of a given disease), requiring decisions about the key entities and relation types to include, the optimal data sources from which to extract information and how best to model the desired knowledge in a KG schema. This manual graph composition process is costly and time-consuming, an issue that can be compounded when KGs are designed as input for ML models, often requiring several iterations to best tune the graph to the predictive task at hand (Bonner et al., 2022a; Ratajczak et al., 2022). A solid understanding of desirable topological properties of KGs that can be leveraged by ML models to make better predictions would allow for a more principled approach to this graph construction process, making it cheaper and faster to generate KGs with higher utility to the practitioner.
Thus, there is a growing interest in understanding how any decisions taken which affect topological properties of a KG can ultimately impact predictions made from them. Amongst other topological properties, patterns such as compositions (Figure 1 (d)), where a relation between two entities is inferred by the presence of a connecting third entity, can be used to link graph structure to model prediction quality. In the biomedical domain, this could be the case when a drug is inferred to treat a disease because they share a common gene connection. Ideally, we would like to discover new drug repurposing opportunities by learning from such patterns in the data. Similarly, the inference pattern (Figure 1 (b)) uses the presence of a certain relation type between an entity pair to infer the presence of a different relation between the same pair, which could, for example, be used to discover novel therapeutic gene-disease connections by aggregating multiple evidence types. Another example is in Protein-Protein Interactions (PPI), where an edge could have a clear directionality if representing information flow in a pathway, or be inherently symmetrical (Figure 1 (a)) if representing simple binary physical interaction.
@@
While the theoretical implications of these topological patterns for KGE models are well understood – in fact, new KGE models are often designed to capture key topological patterns – their impact on the predictive performance in real-world Knowledge Graph Completion tasks remains an open question. Recent work has started to investigate this (Ali et al., 2021; Jin et al., 2023; Teneva & Hruschka, 2023) but has primarily focussed on the properties of relation types rather than individual triples. This approach can help to understand which topological patterns can be captured by the relation embedding when comparing or proposing new KGE models. However, when investigating the impact of the graph structure on KGE model performance, this relation-level approach might neglect critical local properties of the graph. Moreover, existing work has focussed on a set of general-domain KGs, which may contain distinct topological patterns compared to biomedical KGs.
Thus, despite the clear need for understanding the relationship between the topology of KGs and model performance specifically in the biomedical domain, there is a gap in the existing literature, which we aim to address in this work by exploring such relationship at an unprecedented level of detail.
Our main contributions are as follows:
@@
  • We provide an in-depth analysis of the topological properties of six public KGs, focussing on the biomedical domain, and compare the corresponding predictive performance of five well-established KGE models. While previous work has limited the investigation to the macro-level of relation types, we are able to detect stronger patterns linking topological properties to predictive accuracy by zooming in at the level of individual triples.
@@
  • To aid future work in this area, we propose a standardized framework to describe KG topology and release a dedicated Python package kg-topology-toolbox [^1] https://github.com/graphcore-research/kg-topology-toolbox offering a wide range of tools to compute topological statistics of any KG. To help bridge the gap between industry and academia, we also make the predictions of the trained models available to the community (Cattaneo et al., 2024) to conduct further analysis.
@@ 2 Related Work
Since the adoption of representation learning for Knowledge Graph Completion, there has been interest in linking predictive performance of models to topological structures in the underlying Knowledge Graphs (Bordes et al., 2013). These structures, taking into account the heterogeneous relation types in KGs, have been split into two categories: edge cardinalities and topological patterns. For the definitions used in this study see section Knowledge Graph Topological Properties. Indeed, it has become common for new models to be designed with the capacity of capturing these patterns (Cao et al., 2021; Chao et al., 2021; Sun et al., 2019), working under the assumption that better modelling them will result in improved predictive performance (Jin et al., 2023).
@@
There has been limited work providing a comprehensive study on the impact of these patterns across different models and datasets, although some studies have started to appear recently. For example, an investigation across 3 common benchmark KGs (FB15k-237, WN18RR and Yago310) and 21 KGE models found that symmetric relations were easier to predict that anti-symmetric and compositions ones (Ali et al., 2021). Another study compared the performance of 7 KGE models on FB15k-237 and WN18RR, focussing on 4 different topological patterns (Jin et al., 2023), finding that a model’s theoretical capacity to capture a certain pattern does not guarantee good predictive performance on it. Taking a different approach, a study showed how adding noise to a graph impacted the ability of TransE to learn the composition pattern (Douglas et al., 2022). However, none of these studies explores the impact of edge patterns on biomedical KGs specifically.
Biomedical KGs can suffer from degree imbalance, which can arise naturally from the biology modelled by the KG (e.g., key hub-genes having a high degree) or artificially (e.g., due to inspection bias where more attention is paid to high degree entities). This degree imbalance can impact downstream models: through perturbation of graph topology, degree alone has been shown to be a strong prior predictor for the existence of edges (Zietz et al., 2024). KGE models are known to be strongly impacted by entity degree, with models often being heavily biased towards highly connected entities (Bonner et al., 2022b; Rossi et al., 2021b, a; Mohamed et al., 2020). However, disentangling the effect of degree from other topological properties when considering model performance is often overlooked by prior studies. One paper showed that as entity degree increases, performance on symmetric relations decreases (Jin et al., 2023). However, it is still an open research question as to how degree obscures other correlations in the performance of models on different topological patterns.
The prevalence of different topological patterns in real-world datasets is also an active area of research. The frequency of topological and cardinality patterns in 29 real-world KGs from a range of domains, including biomedical was investigated in (Teneva & Hruschka, 2023). The authors observed that, under their chosen pattern definitions, many relations in biomedical KGs are predominantly many-to-many and that there was a lack of composition relations. Other work investigated the prevalence of symmetric and composition relations in 3 common KG benchmark datasets, finding that the majority of relations were anti-symmetric and that composition relations were rarer (Ali et al., 2021).
@@ 3 Knowledge Graph Topological Properties
3.1 Topological Patterns
In terms of describing KG topology, we focus on both the cardinality of the head and tail of triples, as well as their edge topological patterns. Specifically, for a triple $(h,r,t)$ in a KG $\mathcal{G}$ we define the head out-degree $\textrm{deg}(h)$ as the cardinality of the set $\{\hat{t}\mid\exists\hat{r}:(h,\hat{r},\hat{t})\in\mathcal{G}\}$ and, analogously, the tail in-degree $\textrm{deg}(t)$ as the cardinality of $\{\hat{h}\mid\exists\hat{r}:(\hat{h},\hat{r},t)\in\mathcal{G}\}$. We further define the head out-degree of same relation type $\textrm{deg}_{r}(h)$ and the tail in-degree of same relation type $\textrm{deg}_{r}(t)$ as the cardinalities of $\{\hat{t}\mid(h,r,\hat{t})\in\mathcal{G}\}$ and $\{\hat{h}\mid(\hat{h},r,t)\in\mathcal{G}\}$, respectively.
The edge cardinality of a triple $(h,r,t)$ is then defined as:
$\textrm{deg}_{r}(h)$ $\textrm{deg}_{r}(t)$ =1=1 >1>1
=1=1 one-to-one (1:1) many-to-one (M:1)
>1>1 one-to-many (1:M) many-to-many (M:M)
@@ 3.1 Topological Patterns
htrrrr
(a) Symmetry
[…]
(b) Inference
[…]
(c) Inverse
[…]
(d) Composition
Figure 1: The four primary edge topological patterns we consider. We show $(h,r,t)$ as the base triple common to all patterns, while the dashed lines with relations $r^{\prime}$, $r_{1}$ or $r_{2}$, are edges that realize the defining feature of the pattern. We use $r^{\prime}$, $r_{1}$ or $r_{2}$ to denote relations distinct from $r$.
Figure 2: Example of edge cardinalities in a KG with two relation types (blue, red). We use the notation $\textrm{deg}_{r}(t)$$\textrm{deg}_{r}(h)$. For example, for the red edge with cardinality 1:M, the head has more than one outgoing edges of the same relation type (red) and the tail has a single red incoming edge.
@@
While the above cardinalities and topological patterns are well defined for individual triples, previous studies (Jin et al., 2023; Teneva & Hruschka, 2023; Bordes et al., 2013; Sun et al., 2019) have considered them as properties of relation types as a whole (e.g., relation $r$ is said to be symmetric if $\forall h,t,\;\displaystyle{(h,r,t)\Rightarrow(t,r,h)}$), despite the fact that a given property might be satisfied only by a fraction of the triples of a certain relation. We find that such aggregations lack proper formalization in the literature, with inconsistencies in how it is performed, and is prone to introducing noise in results (see Effect of Topological Properties on Model Accuracy). Thus, we consider all the above defined properties at the individual triple level.
3.2 Topological Patterns in Biomedical KGs
To demonstrate how topological patterns can be observed in biomedical KGs, we consider examples from Hetionet.
Symmetric edges: Hetionet contains only a small number of symmetric edges (see Table 1), with all examples being of the Gene-regulates-Gene relation type. In this instance, the symmetric edges represent that knockdown/overexpression of one gene significantly dysregulated the other (Himmelstein et al., 2017). The creation of symmetric edges when the underlying data is not inherently directional (e.g., physical binding interactions between proteins) is a composition design decision that can, depending upon the scoring function chosen, be used when a KGE model shouldn’t learn anything about the directionality of the edge.
Inference edges: Hetionet has several relation types that contain inference edges, for example for Gene-Disease links, where the Disease-upregulates-Gene or Disease-downregulates-Gene relation types are inference edges with Disease-associates-Gene. Inference edges appear when there are logically or semantically distinct relationships between the same pair of entities. They can also arise when the relationship is captured in more than one underlying data source. Here, there is logically an association between a given disease and gene if an impact on gene expression has been measured. The relation types are also drawn from different sources (associates from sources like DisGeNet and GWAS Catalog, up/downregulates from a differential expression analysis (Himmelstein et al., 2017)).
Inverse edges: Inverse edges are rare in Hetionet, with all examples arising from combinations of different Gene-Gene relation types (a pair of genes being linked in one direction by regulation edge and in the other by an interaction edge). These examples may not be intentionally inverse and could be a product of the data integration process. Hypothetical inverse examples (Drug-treats-Disease and Disease-treatedby-Drug for example) could be added to a KG to enable simpler querying by an end user however.
Figure 3: Counts of composition patterns for key relation types within Hetionet and PrimeKG. Values above the bars indicate the number of distinct pairs of relation types in the composition.
Composition edges: Hetionet contains many relation types that are part of composition patterns (for example Compound-binds-Gene being part of the composition pattern Compound-treats-Disease and Disease-associates-Gene), as they naturally arise in biomedical data. To illustrate this further, Figure 3 shows the counts of composition patterns connecting pairs of specific entity types, where we include PrimeKG in addition to Hetionet for comparison. These counts have been stratified for each edge by the type of the intermediate entity $n$ in the composition pattern, and we show the number of unique patterns containing that entity type. This figure shows the abundance and diversity of this pattern across both datasets, where the most abundant composition edge is made up of three genes.
While all these four topological patterns are present in Hetionet, it is worth noting that some of the patterns are inherent to the nature of Biomedical KGs, while others may arise due to construction decisions taken during the creation of the KG.
@@ 4 Experimental Setup
We investigate five public biomedical KGs: Hetionet (Himmelstein et al., 2017), PrimeKG [^2] In the original PrimeKG graph, for every triple $(h,r,t)$ the reverse triple $(t,r,h)$ is present as well. We prepocessed PrimeKG to remove these reverse edges. (Chandak et al., 2023), PharmKG (Zheng et al., 2021), OpenBioLink (Breit et al., 2020), PharMeBINet (Königs et al., 2022) (detailed in Table A.1). These datasets are among the most common biomedical KGs used for benchmarking in the literature (Teneva & Hruschka, 2023) and were chosen with the aim of encompassing a wide variety of sizes and topological patterns. We also include the trivia KG FB15k-237 (Toutanova & Chen, 2015) as a baseline, to detect any results unique to the biomedical domain. On all KGs we train five of the most popular KGE models: TransE (Bordes et al., 2013), DistMult (Yang et al., 2015), RotatE (Sun et al., 2019), TripleRE (Yu et al., 2022) and ConvE (Dettmers et al., 2018). Note that ConvE – in addition to entity and relation embeddings – also learns the parameters of a convolutional network, used in the scoring function (details in Table A.2). Whilst being all well-established among practitioners (Hu et al., 2020), these models differ in complexity and in the relational patterns they can capture (Table A.2). In particular, TransE and DistMult model the interaction between head entity, relation type and tail entity in quite simple terms. Nevertheless, despite missing certain topological patterns, they remain strong baselines. On the other hand, RotatE, TripleRE and ConvE are more recent approaches able to capture all four investigated topological patterns. Notice however that the theoretical capability of a scoring function to model a particular edge topological pattern is not per se a guarantee of stronger predictive performance on such edges (Jin et al., 2023). Our experiments are designed to quantify this impact.